# Ernesto L. Estevanell-Valladares

### v1

@@ Abstract
Experts in machine learning leverage domain knowledge to navigate decisions in model selection, hyperparameter optimisation, and resource allocation. This is particularly critical for fine-tuning language models (LMs), where repeated trials incur substantial computational overhead and environmental impact. However, no existing automated framework simultaneously tackles the entire model selection and HPO task for resource-efficient LM fine-tuning. We introduce XAutoLM, a meta-learning–augmented AutoML framework that reuses past experiences to optimize discriminative and generative LM fine-tuning pipelines efficiently. XAutoLM learns from stored successes and failures by extracting task- and system-level meta-features to bias its sampling toward fruitful configurations and away from costly dead ends. On four text classification and two question-answering benchmarks, XAutoLM surpasses zero-shot optimiser’s peak F1 on five of six tasks, cuts mean evaluation time by up to 4.5×, reduces error ratios by up to sevenfold, and uncovers up to 50% more pipelines above the zero-shot Pareto front. In contrast, simpler memory-based baselines suffer negative transfer. We release XAutoLM and our experience store to catalyze resource-efficient, Green AI fine-tuning in the NLP community.
XAutoLM: Efficient Fine-Tuning of Language Models via Meta-Learning and AutoMLthanks: Work in Progress. Under review at EMNLP 2025. This is not the final version.
Ernesto L. Estevanell-Valladares University of Alicante University of Havana ernesto.estevanell@ua.es          Suilan Estévez-Velarde University of Havana sestevez@matcom.uh.cu
Yoan Gutiérrez University of Alicante          Andrés Montoyo University of Alicante          Ruslan Mitkov University of Alicante Lancaster University
@@ 1 Introduction
Fine-tuning large language models (LMs) has become indispensable across NLP applications, yet even “small” models such as BERT (Devlin et al., 2018) or T5 (Raffel et al., 2020) incur substantial computational cost and carbon emissions (Wang et al., 2023b; Schwartz et al., 2020). Rather than exhaustively evaluating every model and hyperparameter combination, human experts draw on domain knowledge to focus on promising regions of this vast design space.
@@
Automated Machine Learning (AutoML) seeks to mimic expert intuition by automating the two core stages of pipeline construction—model selection (MS) and hyperparameter optimisation (HPO)—into a unified search loop (Hutter et al., 2019). AutoML techniques have matured in areas such as tabular and vision tasks Hutter et al. (2019), showing competitive performance against human experts Estevez-Velarde et al. (2020). However, the joint MS+HPO pipeline for language models presents an ample, mixed discrete-continuous search space whose repeated evaluations are prohibitively costly (Wang et al., 2023b), thus posing a significant challenge for automation. While several recent efforts address HPO for LMs in isolation (Mallik et al., 2024), surveys highlight the underdevelopment of full-pipeline AutoML in NLP (Tornede et al., 2023), and no framework systematically unifies model selection and HPO under tight compute and Green AI constraints.
@@
To address these shortcomings, we propose XAutoLM, the first AutoML framework that unifies model selection and HPO for language model fine-tuning via meta-learning. Rather than initiating a zero-shot search, XAutoLM retrieves a repository of past LM-pipeline evaluations—each annotated with task- and system-level meta-features—to construct an experience-aware prior that biases the search space toward historically fruitful configurations and away from costly failures. By warm-starting the search in this manner, XAutoLM achieves substantial efficiency gains without sacrificing performance. Empirically, across four text classification and two question-answering benchmarks (Table 56), we report reductions in mean evaluation time by up to 4.5×, cuts error ratios by up to sevenfold, and uncovers up to 50% more pipelines dominating the zero-shot Pareto front. All code and our experience store are released [^1] https://anonymous.4open.science/r/XAutoLLM-A010 to accelerate sustainable, reproducible LM fine-tuning in the NLP community.
We summarise the main contributions of our paper as follows:
@@
  • We propose XAutoLM, the first unified, meta-learning–augmented AutoML framework that integrates MSand HPO for discriminative and generative LM fine-tuning.
  • We design an extensible, task- and model-agnostic experience-aware prior mechanism that encodes past LM pipeline successes and failures via task- and system-level meta-features to warm-start the search.
  • We validate XAutoLM on six diverse benchmarks, demonstrating up to 4.5× reduction in search time, sevenfold error-rate reduction, and significant Pareto front gains over zero-shot and naive memory priors.
@@ 2 Related Work
We observe that there are more LLMs for AutoML systems than vice versa, likely due to the proliferation of prompt engineering and increased access to open-source language models. For instance, Zhou et al. (2022) developed the Automatic Prompt Engineer (APE) system, which achieved performance competitive with human-generated instructions. In contrast, systems such as GL-Agent Wei et al. (2023), AutoM3L Luo et al. (2024) and GizaML Sayed et al. (2024) integrate language models into their optimisation strategies to produce graph learning pipelines, highly capable multi-modal ML pipelines, and time-series forecasting pipelines, respectively.
@@
Often, the choice of model is as, if not more, critical than the hyperparameter configuration used to produce responses. We found that AutoGOAL Estevanell-Valladares et al. (2024) optimises pipelines by balancing efficiency and performance metrics, taking into account both model selection and HPO, but only supports LMs for inference. All other AutoML for LLMs systems we surveyed, such as EcoOptiGen Wang et al. (2023a) and PriorBand Mallik et al. (2024), focus solely on HPO.
Nonetheless, we find no single framework that simultaneously addresses model selection and hyperparameter optimisation for LM fine-tuning, primarily when resource limitations exist.
@@ 3 Proposal
We introduce XAutoLM, the first AutoML framework that unifies model selection and hyperparameter optimisation for both discriminative and generative language model fine-tuning. Our pipelines are composed of (i) a base LM from a curated pool of encoders and generators (Table 2), (ii) one of three fine-tuning strategies—full, partial, or LoRA (Hu et al., 2021), and (iii) a hyperparameter configuration. XAutoLM jointly explores this mixed search space by reusing past experiences—e.g., “LoRA-tuned DistilBERT achieved high macro-F1 on SST-2 under low VRAM”—to steer the optimiser toward high-utility regions and away from error-prone configurations. This holistic reuse enables XAutoLM to discover strong fine-tuning pipelines under tight compute budgets.
@@
XAutoLM builds on AutoGOAL’s [^3] Open-source available at: https://github.com/autogoal/autogoal, licensed without restriction. probabilistic optimiser (Estevez-Velarde et al., 2020). The optimiser represents every valid LM pipeline $c$ as a point in a mixed search space that combines discrete choices (e.g. fine-tuning method, model, tokeniser) with continuous hyperparameters (e.g. learning rate, dropout). It maintains a probability distribution $P(c\!\mid\!\theta)$ over that space. It repeats a simple sample–evaluate–update loop: (1) sample a batch of pipelines from $P(c\!\mid\!\theta)$; (2) evaluate them on the target task; and (3) update $P(c\!\mid\!\theta)$ so that high-performing pipelines gain probability mass while under-performing and failures lose it. AutoGOAL always initialises this distribution uniformly, meaning every pipeline—adequate or not—is equally likely at the first generation.
@@ 3.1 Process Overview
XAutoLM replaces this uniform cold start with an experience-aware prior that follows a structured meta-learning process. Initially, the framework retrieves relevant historical evaluations (experiences) from a centralized repository (§ 3.2). Then, it computes detailed task and system meta-features (§ 3.2.1) to characterize the complexity and available resources for the present optimisation task. Leveraging this information, XAutoLM probabilistically adjusts the AutoML search space (§ 3.3), focusing on historically successful configurations and reducing exploration of previously unsuccessful paths. Once configured, the AutoML optimisation starts, fine-tuning pipelines are evaluated, and their outcomes—both successful and unsuccessful—are recorded back into the experience repository, to be used in future runs.
@@ 3.2 Experience Store
[…]
We label an experience positive if all fitness metrics are valid and negative otherwise, usually due to errors occurring during evaluation (out-of-memory, timeout, etc.). Both types are essential: positives pull the search toward fruitful regions, and negatives push it away from costly dead-ends (§ 3.3).
@@ 3.2.1 Meta-Features
We design two complementary meta-feature templates according to the nature of the output space of a task. When the output is drawn from a closed label set—as in text classification or sequence labelling—dataset difficulty is dominated by class imbalance and document-length variation. Conversely, tasks whose output is an open text sequence (question answering, summarisation, translation) demand features that capture the relationship between the input prompt and the target text. Table 3 lists the core features for each template; the same templates can be reused for other label-based or free-form generation tasks with minimal adaptation.
@@
Experiences also hold the hardware characteristics (e.g., number of CPU cores, RAM size) to help condition experiences utility within resource constraints. For instance, while Llama 3.1 70B may yield superior results to smaller alternatives, systems with low VRAM cannot utilise its power.
XAutoLM constructs a holistic representation of each optimisation scenario by combining task-specific and system-level meta-features, enabling robust similarity assessments across diverse contexts.
3.3 Warm-Start optimisation
XAutoLM maintains a probabilistic model $P(c\mid\theta)$ Estevez-Velarde et al. (2020) over pipeline configurations $c$. When a new task $T$ arrives, we retrieve a set of past experiences =$\mathcal{E}=\{e_{1},\dots,e_{n}\}$ and update the model in two sweeps—one for positive experiences, one for negatives:
@@ 4 Experimentation
The second experiment focuses on question answering. We select SQuAD 1.1 Rajpurkar et al. (2016) and DROP Dua et al. (2019) because they share the same input modality yet differ sharply in answer type—extractive spans versus multi-step numerical reasoning—making them a challenging test-bed for generative pipelines. For both studies, experiences are only exchanged among tasks of the same family; classification traces are invisible to QA runs and vice-versa. This constraint ensures that the reported gains stem from task-relevant meta-knowledge rather than accidental data leakage.
@@
Every run is compared against Zero-Shot AutoGOAL, the original optimiser with a uniform sampling distribution; in this setting, the update rules of Eqs. (1)–(9) are never triggered.
In the text-classification study, we include a naive kNN-50 memory baseline. For every target task, we assemble a query vector that concatenates (a) the task meta-features, (b) the current system profile, and (c) the best metric values observed across all stored traces—this encourages the search to drift toward high-performing regions. Distances to positive traces are computed on the full feature+metric space, whereas distances to negative traces ignore metrics (errors lack valid scores). The $k$ nearest positives and $k$ nearest negatives are selected; all receive the same fixed learning rate $\alpha_{i}^{\pm}=1/k$. Setting $u_{i}\!=\!1$ and $\beta\!=\!0$ in Eqs. (3)–(4) reduces our framework to this simple neighbour rule. For question answering the repository contains only between 5 and 10 positive traces per source task, making a neighbour count unreliable; therefore Zero-Shot remains the sole baseline in that study.
@@
Throughout the paper, a pipeline configuration denotes the concrete LM, fine-tuning recipe and hyper-parameters that AutoGOAL evaluates. A warm-start prior (WS Prior) instead refers to the hyper-parameters that govern the probability update equations—distance type, utility scheme, decay factor $\beta_{\text{scale}}$, and the pull limits $(k_{\text{pos}},k_{\text{neg}})$.
For every target task we enumerate meaningful combinations of these prior parameters, yielding roughly a hundred and eighty candidates. Before any optimisation begins we measure the total-variation (TV) distance between each candidate prior and the uniform distribution, sort the list, and slice it into three bins that we label low, moderate and high bias. In classification, we evaluate the median TV and the maximum TV Priors from every bin (six variants); in QA, we keep only the maximum TV Prior from each bin (three variants) to limit computing. Full probability plots and the identifiers of the selected priors appear in Appendix B.
Optimisation protocol.
In every AutoML run, every discovered LM pipeline receives one and a half GPU hours in Text Classification and two hours in QA. Objectives are $\langle F1_{\text{macro}},ET\rangle$ for classification and $\langle F1,ET\rangle$ for QA, where $ET$ is wall-clock evaluation time in seconds. All searches share a fixed random seed of 42; therefore, differences arise solely from the chosen warm-start prior.
Experience repositories.
All warm-started AutoML runs used an experience store (§ 3.2) generated by the previous 48 hours Zero-shot runs on each task. Table 4 lists the experiences produced and the total experiences available for warms-start priors per task. The warm-start mechanism accesses only experiences originating from other tasks in the same study for a clean cross-task evaluation.
@@ 4.1 Text Classification Results
Table 5 summarises the effect of warm-start (WS) priors on the four classification benchmarks. We report a diverse set of metrics of both performance and efficiency: max and mean $F_{1}$ scores reflect peak and average classification quality, mean evaluation time (ET) captures resource cost, error ratio indicates robustness, and hypervolume (HV) measures the coverage of the Pareto front in objective space (Zitzler and Thiele, 1998). Zero-shot performance values correspond to the first 24 hours of their 48 hours initial execution.
Across datasets, WS priors either match or surpass the best Zero-shot $F1_{\text{m}}$ while systematically improving efficiency. On LIAR, a High prior lifts peak $F1_{\text{m}}$ from 0.24 to 0.26, cuts the mean $ET$ by a factor of 3.5, and lowers the error ratio by a sevenfold. A similar pattern emerges on MELD, where High drives the error ratio from 0.77 to 0.10 and reduces mean $ET$ 4.5×, while keeping $F1_{\text{m}}$ above the baseline. On SST-2, the Zero-shot baseline generated the highest $F1_{\text{m}}$ and lowest $ET$ out of all variants. However almost all WS priors reduced Error Ratio and Low (Max) in particular showed a sensible reduction in mean $ET$ while maintaining peak $F1_{\text{m}}$. On AG News, all WS variants improve max $F1_{\text{m}}$ while several improve $ET$, hypervolume and Error Ratio, showing that better performance–time trade-offs are discoverable even in large-scale settings.
The naïve kNN-50 baseline, although in SST-2 case attains large HV values, degrades performance on three datasets, notably obtaining the worst results out of all priors in AG NEWS (0.90 → 0.67 $F1_{\text{m}}$) and MELD (0.41 → 0.37 $F1_{\text{m}}$).
@@ 4.2 Question Answering Results
Table 6 reports results on the generative SQuAD 1.1 and DROP datasets. Knowledge reused from a single related task already yields substantial gains. For SQuAD, WS priors outperform the baseline in almost all metrics. The High-Max prior, in particular, raises $F1$ from 0.34 to 0.89 while shrinking mean ET from 4081s to 1337s (-3×). On DROP, the Low prior illustrates negative transfer, yet both Moderate and High priors outperform Zero-shot on every metric—peak $F1$ improves slightly (0.39→0.40) and mean ET falls by 47%. These outcomes confirm that cross-task meta-knowledge generalises beyond classification and that the adaptive pull/push schedule mitigates catastrophic transfers.

### v2

# XAutoLM: Efficient Fine-Tuning of Language Models via Meta-Learning and AutoML

@@ Abstract
Experts in machine learning leverage domain knowledge to navigate decisions in model selection, hyperparameter optimization, and resource allocation. This is particularly critical for fine-tuning language models (LMs), where repeated trials incur substantial computational overhead and environmental impact. However, no existing automated framework simultaneously tackles the entire model selection and hyperparameter optimization (HPO) task for resource-efficient LM fine-tuning. We introduce XAutoLM, a meta-learning–augmented AutoML framework that reuses past experiences to optimize discriminative and generative LM fine-tuning pipelines efficiently. XAutoLM learns from stored successes and failures by extracting task- and system-level meta-features to bias its sampling toward valuable configurations and away from costly dead ends. On four text classification and two question-answering benchmarks, XAutoLM surpasses zero-shot optimizer’s peak $F1$ on five of six tasks, cuts mean evaluation time of pipelines by up to 4.5×, reduces search error ratios by up to sevenfold, and uncovers up to 50% more pipelines above the zero-shot Pareto front. In contrast, simpler memory-based baselines suffer negative transfer. We release XAutoLM and our experience store to catalyse resource-efficient, Green AI fine-tuning in the NLP community.
XAutoLM: Efficient Fine-Tuning of Language Models via Meta-Learning and AutoML
Ernesto L. Estevanell-Valladares1,2, Suilan Estévez-Velarde2, Yoan Gutiérrez1, Andrés Montoyo1, Ruslan Mitkov3, 1University of Alicante, 2University of Havana, 3University of Lancaster, ernesto.estevanell@ua.es
@@ 1 Introduction
Fine-tuning large language models (LLMs) has become indispensable across natural language processing (NLP) applications, yet even “small” models such as BERT (Devlin et al., 2018) or T5 (Raffel et al., 2020) incur substantial computational cost and carbon emissions (Wang et al., 2023b; Schwartz et al., 2020). Rather than exhaustively evaluating every model and hyperparameter combination, human experts draw on domain knowledge to focus on promising regions of this vast design space.
@@
Automated Machine Learning (AutoML) seeks to mimic expert intuition by automating the two core stages of pipeline construction, model selection (MS) and hyperparameter optimization (HPO), into a unified search loop (Hutter et al., 2019). AutoML techniques have matured in areas such as tabular and vision tasks Hutter et al. (2019), showing competitive performance against human experts Estevez-Velarde et al. (2020). However, the joint MS+HPO pipeline for language models presents an ample, mixed discrete-continuous search space whose repeated evaluations are prohibitively costly (Wang et al., 2023b), thus posing a significant challenge for automation. While several recent efforts address HPO for LMs in isolation (Mallik et al., 2024), surveys highlight the underdevelopment of full-pipeline AutoML in NLP (Tornede et al., 2023), and no framework systematically unifies model selection and HPO under tight compute and Green AI constraints.
@@
To address these shortcomings, we present XAutoLM, an AutoML framework that unifies model selection and hyperparameter optimization for LM fine-tuning via meta-learning. XAutoLM constructs an experience-aware prior from a repository of past pipeline evaluations annotated with task- and system-level meta-features which steers the search toward historically promising and away from infeasible configurations. Empirically, across four classification and two question-answering benchmarks, our method yields pipelines with stronger performance-time trade-offs than zero-shot or naive baselines under identical wall-clock budgets (Tables 56). We release the code and the full experience store [^1] https://github.com/EEstevanell/XAutoLM to support sustainable, reproducible LM fine-tuning in the NLP community.
We summarize our main contributions as follows:
@@
  • A unified, meta-learning–augmented AutoML framework that integrates both model selection and hyperparameter optimisation for discriminative and generative LM fine-tuning.
  • An extensible, task- and model-agnostic experience-aware prior that conditions the search on task and system meta-features and explicitly leverages negative traces to avoid costly dead ends.
  • A comprehensive evaluation on six benchmarks showing consistent gains in $F_{1}$, mean pipeline evaluation time, and error ratio, and stronger Pareto fronts than zero-shot and naive memory baselines (see Section 4; Tables 56).
We next review related work (Section 2), present XAutoLM (Section 3), and report the experimental setup and results (Section 4), followed by analysis (Section 5) and, finally, conclusions and limitations (Sections 67).
@@ 2 Related Work
We observe that there are more LLMs for AutoML systems than vice versa, likely due to the proliferation of prompt engineering and increased access to open-source LMs. For instance, Zhou et al. (2022) developed the Automatic Prompt Engineer (APE) system, which achieved performance competitive with human-generated instructions. In contrast, systems such as GL-Agent Wei et al. (2023), AutoM3L Luo et al. (2024) and GizaML Sayed et al. (2024) integrate language models into their optimization strategies to produce graph learning pipelines, highly capable multi-modal ML pipelines, and time-series forecasting pipelines, respectively.
@@
Often, the choice of model is as, if not more, critical than the hyperparameter configuration used to produce responses. We found that AutoGOAL Estevanell-Valladares et al. (2024) optimizes pipelines by balancing efficiency and performance metrics, taking into account both model selection and HPO, but only supports LMs for inference. All other AutoML for LLMs systems we surveyed, such as EcoOptiGen Wang et al. (2023a) and PriorBand Mallik et al. (2024), focus solely on HPO.
Nonetheless, we find no single framework that simultaneously addresses model selection and hyperparameter optimization for LM fine-tuning, particularly when resource limitations exist.
@@ 3 Proposal
We introduce XAutoLM, the first AutoML framework that unifies model selection and hyperparameter optimisation for both discriminative and generative language model fine-tuning. Our pipelines are composed of (i) a base LM from a curated pool of encoders and generators (Table 2), (ii) one of three fine-tuning strategies; full, partial, or LoRA (Hu et al., 2021), and (iii) a hyperparameter configuration. XAutoLM jointly explores this mixed search space by reusing past experiences, e.g., “LoRA-tuned DistilBERT achieved high macro-F1 on SST-2 under low VRAM”, to steer the optimizer toward high-utility regions and away from error-prone configurations. This holistic reuse enables XAutoLM to discover strong fine-tuning pipelines under tight compute budgets.
@@
XAutoLM builds on AutoGOAL’s [^3] Open-source available at: https://github.com/autogoal/autogoal, licensed without restriction. probabilistic optimizer (Estevez-Velarde et al., 2020). The optimizer represents every valid LM pipeline $c$ as a point in a mixed search space that combines discrete choices (e.g. fine-tuning method, model, tokenizer) with continuous hyperparameters (e.g. learning rate, dropout). It maintains a probability distribution $P(c\!\mid\!\theta)$ over that space. It repeats a simple sample–evaluate–update loop: (1) sample a batch of pipelines from $P(c\!\mid\!\theta)$; (2) evaluate them on the target task; and (3) update $P(c\!\mid\!\theta)$ so that high-performing pipelines gain probability mass while under-performing and failures lose it. AutoGOAL always initializes this distribution uniformly, meaning every pipeline, adequate or not, is equally likely at the first generation.
@@ 3.1 Process Overview
XAutoLM replaces this uniform cold start with an experience-aware prior that follows a structured meta-learning process. Initially, the framework retrieves relevant historical evaluations (experiences) from a centralized repository (Section  3.2). Then, it computes detailed task and system meta-features (Section  3.2.1) to characterize the complexity and available resources for the present optimisation task. Leveraging this information, XAutoLM probabilistically adjusts the AutoML search space (Section  3.3), focusing on historically successful configurations and reducing exploration of previously unsuccessful paths. Once configured, the AutoML optimisation starts, fine-tuning pipelines are evaluated, and their outcomes, both successful and unsuccessful, are recorded back into the experience repository, to be used in future runs.
@@ 3.2 Experience Store
Our system learns from a growing repository of experiences; past pipeline evaluations that capture every factor influencing performance. Formally, an experience is a 4-tuple $e=\langle c,\,\mathbf{m},\,t,\,s\rangle$ where $c$ is the complete pipeline configuration, $\mathbf{m}$ the vector of recorded metrics (e.g. $\mathrm{F1}$, ROUGE, evaluation time), $t$ a task meta-feature vector, and $s$ straightforward system descriptors such as CPU cores, RAM, and GPU memory.
We label an experience positive if all fitness metrics are valid and negative otherwise, usually due to errors occurring during evaluation (out-of-memory, timeout, etc.). Both types are essential: positives pull the search toward valuable regions, and negatives push it away from costly dead-ends (Section  3.3).
@@ 3.2.1 Meta-Features
We design two complementary meta-feature templates according to the nature of the output space of a task. When the output is drawn from a closed label set, as in text classification or sequence labelling, dataset difficulty is dominated by class imbalance and document-length variation. Conversely, tasks whose output is an open text sequence (question answering, summarisation, translation) demand features that capture the relationship between the input prompt and the target text. Table 3 lists the core features for each template; the same templates can be reused for other label-based or free-form generation tasks with minimal adaptation.
@@
Experiences record a minimal hardware profile in s (CPU cores, CPU frequency, total RAM, GPU VRAM) so similarity and feasibility reflect both task and system characteristics. For instance, while Llama 3.1 70B may yield superior results to smaller alternatives, systems with low VRAM cannot utilize its power.
XAutoLM constructs a holistic representation of each optimization scenario by combining task-specific and system-level meta-features, enabling robust similarity assessments across diverse contexts.
3.3 Warm-Start optimization
XAutoLM maintains a probabilistic model $P(c\mid\theta)$ Estevez-Velarde et al. (2020) over pipeline configurations $c$. When a new task $T$ arrives, we retrieve a set of past experiences $\mathcal{E}=\{e_{1},\dots,e_{n}\}$ and update the model in two sweeps; one for positive experiences, one for negatives:
@@ 4 Experimentation
The second experiment focuses on question answering. We select SQuAD 1.1 Rajpurkar et al. (2016) and DROP Dua et al. (2019) because they share the same input modality yet differ sharply in answer type, extractive spans versus multi-step numerical reasoning, making them a challenging test-bed for generative pipelines. For both studies, experiences are only exchanged among tasks of the same family; classification traces are invisible to QA runs and vice-versa. This constraint ensures that the reported gains stem from task-relevant meta-knowledge rather than accidental data leakage.
@@
Every run is compared against Zero-Shot AutoGOAL, the original optimizer with a uniform sampling distribution; in this setting, the update rules of equations (1)–(9) are never triggered.
In the text classification study, we include a naive kNN-50 memory baseline for comparing against a naive experience retrieval method. For every target task, we assemble a query vector that concatenates (a) the task meta-features, (b) the current system profile, and (c) the best metric values observed across all stored traces; this encourages the search to drift toward high-performing regions. Distances to positive traces are computed on the full feature+metric space, whereas distances to negative traces ignore metrics (errors lack valid scores). The $k$ nearest positives and $k$ nearest negatives are selected; all receive the same fixed learning rate $\alpha_{i}^{\pm}=1/k$. Setting $u_{i}\!=\!1$ and $\beta\!=\!0$ in equations (3)–(4) reduces our framework to this simple neighbour rule. For question answering the repository contains only between 5 and 10 positive traces per source task, making a neighbour count unreliable; therefore Zero-Shot remains the sole baseline in that study.
@@
[…]
For each task, we enumerate $\approx 180$ WS-prior parameterizations. For a given candidate prior to a task, we apply it with the fixed experience store (leaving the experience for the current task out) to obtain the induced sampling distribution $p$ over fine-tuning methods on that task. We then compute the total-variation (TV) distance between this induced marginal and the uniform distribution over the same method set. We rank candidates by TV and split them into three data-driven strata (low | moderate | high bias) at prominent TV gaps ($\approx 2\times$). In classification, we select per strata the median-TV and max-TV priors (six priors total). In QA, we select only the max-TV prior per strata (three priors) to respect the compute budget. Full probability plots of the induced method distributions and the selected prior identifiers are provided in Appendix B.
Execution protocol.
For each task, we first ran the Zero-shot configuration for 48 hours to populate the experience store. Table 4 reports the positive/negative traces generated by this baseline run on each task. We then executed the kNN-50 baseline and all WS-prior variants for 24 hours of wall-clock time each. The warm-start mechanism accesses only experiences originating from other tasks within the same study (clean cross-task transfer; see Table 4). For fairness in reporting, Zero-shot metrics are computed from the first 24 hours of their 48 hours runs, matching the wall-time allocated to WS-priors and kNN-50. This protocol isolates whether experience improves both effectiveness and efficiency under the same time budget.
@@
In every AutoML run, each discovered LM pipeline has up to 1.5 GPU-hours in Text Classification and 2 GPU-hours in QA for evaluation. Objectives are $\langle F1_{\text{macro}},ET\rangle$ for classification and $\langle F1,ET\rangle$ for QA, where $ET$ is the wall-clock evaluation time of a pipeline (in seconds). All searches share a fixed random seed ($42$) and the same hardware; therefore, differences arise solely from the chosen warm-start prior.
@@ 4.1 Text Classification Results
Table 5 summarizes the effect of WS-priors on the four classification benchmarks. We report both performance and efficiency: max and mean $F1_{\text{macro}}$ reflect peak and average classification quality; mean evaluation time (ET) captures resource cost; the error ratio indicates the share of failed pipeline evaluations; and hypervolume (HV) measures Pareto-front coverage in objective space (Zitzler and Thiele, 1998). Mean ET is averaged over successfully completed pipeline evaluations only (i.e., runs that return valid fitness metrics); failed evaluations (e.g., out-of-memory, timeouts, runtime errors) are excluded from ET and are accounted for by the error ratio. All methods are run under the same 24 hours single-GPU budget (cf. Execution protocol), so ET differences reflect pipeline runtime rather than total search compute.
@@
Across datasets, WS priors either match or surpass the best Zero-shot $F1_{\text{m}}$ while systematically improving efficiency. On LIAR, a High prior lifts peak $F1_{\text{m}}$ from 0.24 to 0.26, cuts the mean $ET$ by a factor of 3.5, and lowers the error ratio by sevenfold. A similar pattern emerges on MELD, where High drives the error ratio from 0.77 to 0.10 and reduces mean $ET$ 4.5×, while keeping $F1_{\text{m}}$ above the baseline. On SST-2, the Zero-shot baseline generated the highest $F1_{\text{m}}$ and lowest $ET$ out of all variants.
Zero‑shot runs exhibit high error ratios across all benchmarks (e.g., 0.73-0.92); the WS priors cut these failure rates dramatically, down to 0.09-0.90. Moreover, non-naive warm-started runs showed a sensible reduction in mean $ET$ while maintaining peak $F1_{\text{m}}$. On AG News, all WS runs improve max $F1_{\text{m}}$ while several improve $ET$, HV and Error Ratio, showing that better performance–time trade-offs are discoverable even in large-scale settings.
The naive kNN-50 baseline, although in SST-2 case attains large HV values, degrades performance on three datasets and notably obtains the worst results out of all priors in AG NEWS (0.90 → 0.67 $F1_{\text{m}}$) and MELD (0.41 → 0.37 $F1_{\text{m}}$).
4.2 Question Answering Results
Table 6 reports results on the generative SQuAD 1.1 and DROP datasets. Knowledge reused from a single related task already yields substantial gains. For SQuAD, WS priors outperform the baseline in almost all metrics. The High-Max prior, in particular, raises $F1$ from 0.34 to 0.89 while shrinking mean ET from 4081s to 1337s (-3×). Similarly to the text classification results, WS priors bring error ratios down from 0.94–0.95 (zero‑shot) to 0.76–0.90.
On DROP, the Low prior illustrates negative transfer, yet both Moderate and High priors outperform Zero-shot on every metric; peak $F1$ improves slightly (0.39→0.40) and mean ET falls by 47%. These outcomes confirm that cross-task meta-knowledge generalizes beyond classification and that the adaptive pull/push schedule mitigates catastrophic transfers.